Crane Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $46 throughout the country to loyal alumni of over 1,000 schools. Crane’s variable costs are 40% of sales; fixed costs are $120,000 per month.

Assume that variable costs increase to 46% of the current sales price and fixed costs increase by $11,100 per month. If Crane were to raise its sales price by 12% to cover these new costs, what would be the new annual breakeven point in sales dollars? (Round sales price to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)
Answer #1

Answer :- Calculation of New annual break even point in sales dollar :-

New Sales Price = $46 *112% = $51.52

New Variable Cost = $46 *46% = $21.16

New Contribution Margin = $51.52 - $21.16 = $30.36

New Contribution Margin Ratio = $30.36 / $51.52 = 58.92857%

Fixed Cost per year = ($120,000 + $11,100) *12 = $1,573,200

Break Even point in sales dollar = Fixed Cost / CM Ratio

Break Even point in sales dollar =$1,573,200 / 58.92857%

Break Even point in sales dollar =$2,669,672.79 OR $2,669,673

If you have any query ask in comment section. If you like the answer plz rate. Thanks